If you stand on a bathroom scale, the spring inside the scale compresses 0.80 mm, and it tells you your weight is 700 N.
Now if you jump on the scale from a height of 1.8 m, what does the scale read at its peak?
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2019-04-01T10:32:40-0400
"kx=W"

"k=\frac{700}{0.0008}=875000 N/m"

"0.5kx^2=Wh"

"x=\sqrt{\frac{2Wh}{k}}=\sqrt{\frac{2(700)(1.8)}{875000}}=0.0537 m."

So,


"F=kx=(875000)(0.0537)=47000N."